OpenStudy (anonymous):

Fan and medal reward Which function is the inverse of f(x)=x^3−6x^2+12x−8?

OpenStudy (anonymous):

f^(-1) (x) = x^(1\/3)+2

OpenStudy (anonymous):

is it \[f^-1(x)=\sqrt[3]{x}+2\] or \[f^-1(x)=\sqrt[3]{x+2}\] @Famous5

OpenStudy (anonymous):

@dmndlife24

OpenStudy (dmndlife24):

f(x)=x^3−6x^2+12x−8 that function is the same as... f(x)=(x-2)^3 Now to find the inverse just interchange the x and y variables and solve for y... (first change f(x) to y) y=(x-2)^3 Interchange... x=(y-2)^3 Now solve for y to find that it is equivalent to \[y = \sqrt[3]{x+2}\]

OpenStudy (dmndlife24):

Lastly, just substitute back in the "f-1(x)"

OpenStudy (anonymous):

thank you so much! @dmndlife24

OpenStudy (dmndlife24):

No problem